OpenStudy (mackenzie2013):

Evaluate the expression: -4 - [5 - (-5)] ÷ |6 - 8| + 4 5 0 -1 -5

OpenStudy (shamim):

\[-4-[5-(-5)]\div |6-8|+4\]

OpenStudy (shamim):

\[-4-[5+5]\div2+4=-4-5+4=-5\]

OpenStudy (jhannybean):

\[\large -4-[5-(-5)] \div |6-8|+4\]\[\large -4-(10) \div 2+4\]\[\large -4-5+4\]\[\large -4-1\]\[\large= -5\]